# Noisy Quantum Learning Theory

### v1

@@ 2.2 Purity testing in noisy quantum experiments
So far our results have established oracle separations between relativized complexity classes. We now turn to more concrete noisy learning tasks, asking how sample complexity scales with coherent quantum memory and access to joint measurements, and whether exponential quantum advantages can survive constant local noise. Our first result in this direction, Theorem 2.4, addresses a canonical task in quantum learning theory and property testing, namely testing the purity of a quantum state, which has been highlighted as one of the few examples exhibiting quantum advantages in both sample and computational complexity [MdW18, BCL20, HKP21, ACQ22, HBC+22, CCHL22, CGY24, GHYZ24]. In the ideal setting, distinguishing an $n$-qubit maximally mixed state from a fixed pure state requires $\Omega(2^{n})$ samples, whereas permitting joint measurements on two copies reduces the task to only $O(1)$ samples and constant computational time. This separation is powered by a quantum subroutine that appears throughout the literature on quantum speedups in learning, property testing, and computation: the multi-qubit SWAP operation [HM13, LMR14, BOW19, CSSC18, MdW18, HBC+22]. Here, we show that this super-exponential quantum speedup is completely degraded by only a single layer of noise.
@@
Our analysis proceeds by bounding the likelihood ratio between the leaf distributions generated by this learning tree when its measurement outcomes arise from the maximally mixed state versus from a fixed, Haar-random pure state. To do so, we focus on a fixed root-to-leaf path, utilizing the tree’s multilinear structure to reinterpret the learning problem. Under our reinterpretation, it suffices to bound the concentration of the output distributions of intermediate nodes in the tree, which we achieve by proving several technical lemmas regarding the noise channel and using the martingale formalism from [CGY24]. Because each node concentrates by a factor exponentially small in the noise rate, every experiment is highly uninformative, and we obtain our lower bound.
@@
While we have shown that assessing the purity of a generic quantum state is fundamentally obstructed by noise, there are natural classes of “noise-robust” states for which purity testing remains feasible. A toy but instructive example arises in tensor-network models of the AdS/CFT correspondence [Mal99], such as the HaPPY code [PYHP15]. In essence, a HaPPY tensor network $T$ defines an isometric encoding $T:\mathcal{H}_{\text{bulk}}\to\mathcal{H}_{\text{bdy}}$ from a collection of “bulk” logical qudits on a finite hyperbolic lattice to “boundary” physical qudits arranged on a one-dimensional ring. In the AdS/CFT interpretation, $\mathcal{H}_{\text{bulk}}$ models a code subspace of low-energy quantum-gravitational degrees of freedom on a spatial slice of AdS, while $\mathcal{H}_{\text{bdy}}$ models the Hilbert space of the dual CFT on the boundary circle. The quantum error-correcting structure of $T$ discretely implements bulk-boundary reconstruction: local bulk operators can be reconstructed on many overlapping boundary regions, and the radial direction of the network plays the role of a renormalization group scale [JE21].
@@
This result should be viewed as a holographic counterpart of our noisy purity-testing lower bound. On the one hand, the HaPPY code converts local boundary erasures into highly suppressed logical errors on the bulk black hole degrees of freedom: as long as $R\gtrsim r+O(\log r)$ and the erasure rate is below threshold, a greedy decoder can approximately recover $\rho_{\mathrm{BH}}$ from the noisy boundary state with failure probability that is exponentially small in $4^{R-r}$. More broadly, it is believed (but not known) that HaPPY-type holographic codes may admit fully fault-tolerant realizations against local noise [HCM+20, JE21, FHMS21]; in our setting we only appeal to their rigorously understood erasure-correction properties. Composing the decoding map with a two-copy SWAP test on the recovered bulk region therefore reproduces, up to small decoding errors, the ideal two-copy purity test and yields a constant-copy quantum protocol. On the other hand, any protocol restricted to single-copy measurements on the boundary reduces, via bulk-boundary isometry and data-processing, to single-copy purity testing on an $L_{r}$-qubit system, where $L_{r}=\Theta(4^{r-1})$ is the number of bulk legs in the excised region. Our general lower bound for noisy single-copy purity testing then implies a sample complexity of order $2^{\Theta(L_{r})}=2^{\exp(\Theta(r))}$, establishing the separation claimed in the theorem.
2.3 Noise-dependent quantum advantage in Pauli shadow tomography
@@
To further understand how noise reshapes quantum learning advantages, we now consider the well-studied problem of Pauli shadow tomography, namely estimating expectation values of (potentially mutually noncommuting) Pauli observables, which in ideal settings exhibits an exponential memory-sample tradeoff [CCHL22]. Any noiseless conventional protocol restricted to single-copy measurements, or to $k<n$ qubits of quantum memory, requires a number of samples exponential in nkn-k, whereas a quantum-enhanced learner with two-copy access (i.e. $n$ ancillary qubits of quantum memory) succeeds with only $O(n)$ samples using Bell-basis measurements; these two settings are depicted in Fig. 3(a) (memoryless or small-memory single-copy experiments) and Fig. 3(b) (architectures with a quantum memory register enabling multi-copy measurements). In Theorem 2.6, we refine this separation in the NBQP setting, quantifying the noise-dependence of lower bounds with and without ancillary quantum memory. Notably, even when $n$ qubits of memory are provided, enabling two-copy measurements, we show that order $(1-\lambda)^{-n}$ samples are necessary. We give a noisy 2-copy algorithm achieving this scaling up to a constant factor in the exponent, establishing a quantum advantage in noisy Pauli tomography that depends polynomially on the noise rate, and negating the exponential speedup of ideal Bell-measurement-based strategies.
@@
In the presence of constant depolarizing noise per qubit, any quantum algorithm without ancillary quantum memory which can identify a Pauli-structured state with high probability requires order $2^{n}f(\lambda)^{n}$ measurements, where $f(\lambda)\in[1,\infty)$ for $\lambda\in[0,1]$. Given an additional $k\leq n$ qubits of quantum memory, $\Omega(2^{n-k}(1-\lambda)^{-n})$ samples are still required. When $k=n$, there exists a quantum-enhanced learning algorithm with access to noisy two-copy measurements solving the Pauli-identification task using $\tilde{O}((1-\lambda)^{4n})$ samples
As an immediate corollary, we find a sample complexity separation between our two-copy algorithm and any single-copy strategy for the same Pauli identification task. This separation depends on the noise rate, and interpolates smoothly between the ideal exponential separation for Pauli tomography and a polynomial (in the noise rate) separation for $O(1)$ local noise. A related separation is presented in Ref. [HFP22]; that work considers diamond-norm gate noise and access to perfect state copies, a weaker noise model than NBQP, and gives a lower bound which considers only noiseless conventional experiments.
Corollary 2.7.
Let $N_{SC}$ be the optimal sample complexity for any single-copy algorithm for the $n$-qubit Pauli identification task from Theorem 2.6, and let $N_{TC}$ be the sample complexity of the two-copy algorithm given in Theorem 2.6. Then $N_{SC}=\Omega\big(N_{TC}^{a(\lambda)}\big)$, where $a(\lambda)$ is a function of noise rate such that when $\lambda=\Theta(1)$ and $n$ is large, $a(\lambda)=\Theta(\lambda^{-1})$, and as $\lambda\rightarrow 0$, $N_{SC}=\exp(\Omega(n))$.
@@
We next define learning trees for each experimental model. For models (i) and (ii), the output of each node of the tree is a classical bitstring. For both, we show that the distribution over bitstrings induced by the results of each experiment, given access to a $\mathds{1}+P$-type state, concentrate around the distribution resulting from experiments which instead query a maximally mixed state. Hence, every experiment is uninformative in distinguishing the two hypotheses, and exponentially many measurements are required to achieve an algorithm output distribution under one hypothesis that is far in total variation from the other (Theorem F.6 and Theorem F.8). The relationship between concentration of node-wise distributions and the final leaf output distribution is given by the martingale formalism developed in [CGY24]. In model (ii), obtaining this bound requires quantifying the correlation between copies of the unknown state in terms of the size of the quantum memory and its lifetime, which we accomplish using a reformulation of the learning model in terms of Matrix Product States (Lemma F.11).
@@
For model (iii), nodes of the tree are joined by the state of the ancillary memory register after every measurement rather than a simple classical bitstring. Thus, we require a stronger approach: bounding the variation in entire root-to-leaf paths in the learning tree, following the approach from [CCHL22]. Using a probabilistic argument, we control the total variation of the leaf output distribution by bounding the number of paths which diverge substantially from the output distribution induced by the maximally mixed input state. To account for depolarizing noise in every copy of the state, we simplify the tensor-network analysis of [CCHL22] via a technical argument which provides a cleaner form of the action of our noise channel. With this approach, we find that the contribution of the noise to the sample complexity decouples from the number of memory qubits; hence, even given $n$ ancillary memory qubits, $\Omega((1-\lambda)^{-n/3})$ samples are still required in the noisy regime. We provide a two-copy algorithm which leverages noisy Bell sampling to match this asympotic lower bound up to constants in the exponent, and give a single-copy classical shadow algorithm for the broader Pauli shadow tomography which accounts for noise by using shallow-circuit unitary ensembles. Up to polynomial factors and constants in the exponents, our upper and lower bounds for the Pauli-identification task exhibit the same exponential dependence on $n$, $k$, and the noise rate $\lambda$, matching the known noiseless bounds from [CCHL22, CGY24] as $\lambda\to 0$ and diverging to infinity as $\lambda\to 1$, where the task becomes information-theoretically impossible.
@@ 3 Outlook
We have studied quantum learning theory in the presence of noise, with a particular focus on noisy quantum learning-enhanced experiments. Our results show that many of the most striking idealized quantum advantages in learning and property testing disappear, or become inapplicable, once interstitial noise and unprotected oracle systems are taken into account. At the same time, we identified settings where advantages survive, often in a weakened but still meaningful form, and clarified how these surviving advantages depend on noise strength, memory resources, and problem structure.
@@
Our work suggests several concrete directions for further development of noisy quantum learning theory. First, progress will require engaging more deeply with the physics of quantum many-body systems, rather than treating oracles as abstract. Many natural systems possess built-in robustness: renormalization-group structure and locality can protect long-range correlations [KK17, FLO22, FLM22, GLL24, LBC25] , and thermalizing dynamics can generate noise-resilient macroscopic observables [KKCG25, BCSB19, CRL10]. It will be important to identify and characterize classes of states, channels, and observables whose relevant features are intrinsically stable under realistic error models, and to turn such structure into quantitatively sharp, physically natural examples of noisy quantum advantage.
@@
Second, most known exponential quantum advantages in learning rest on highly entangled, global measurements, such as large-scale SWAP tests or Bell-basis measurements [ACQ22, HFP22, CCHL22], that are simultaneously fragile to noise and misaligned with the local-control constraints of real devices which can couple to and manipulate experimental systems. This points toward a theory of noisy quantum learning under restricted resources, in which allowed operations may be shallow, geometrically local, few-qubit, or constrained to a small number of probe systems. Quantum probe tomography [CCH25] is an example of this philosophy, wherein a small number of probes interacting locally with a large system can still extract global information under noise. It will be important to understand systematically which restricted-access models admit robust (even if only polynomial) advantages over conventional experiments, and how those advantages trade off against architectural constraints.
@@
There is a close connection between noisy quantum learning and quantum sensing. Many sensing protocols can be viewed as learning problems about Hamiltonian parameters or state observables under stringent access and noise constraints [ZZPJ18, HTFS23, HMG+25]. Some ambitious proposals, such as quantum computation-enhanced sensing based on deep coherent control [HTFS23, HMG+25] or Grover-type oracles [AMC+25], lose their asymptotic advantage in the NBQP noise model. This raises a quantitative question rather than a purely asymptotic one: how does the achievable quantum advantage degrade as a function of the noise rate, circuit depth, and available memory? Corollary 2.7, together with related work such as Ref. [HFP22], suggest that in realistic noise regimes one should expect noise-dependent polynomial improvements rather than noise-independent exponential ones, but those polynomial advantages can still be practically meaningful.

### v2

@@ 2.2 Purity testing in noisy quantum experiments
So far our results have established oracle separations between relativized complexity classes. We now turn to more concrete noisy learning tasks, asking how sample complexity scales with coherent quantum memory and access to joint measurements, and whether exponential quantum advantages can survive constant local noise. Our first result in this direction, Theorem 2.4, addresses a canonical task in quantum learning theory and property testing, namely testing the purity of a quantum state, which has been highlighted as one of the few examples exhibiting quantum advantages in both sample and computational complexity [montanaro2018surveyquantumpropertytesting, bubeck2020entanglementnecessaryoptimalquantum, Huang_2021, Aharonov_Cotler_Qi_2022, Huang_2022, chen2021exponentialseparationslearningquantum, weiyuan_paulis, gong2024samplecomplexitypurityinner]. In the ideal setting, distinguishing an $n$-qubit maximally mixed state from a fixed pure state requires $\Omega(2^{n})$ samples, whereas permitting joint measurements on two copies reduces the task to only $O(1)$ samples and constant computational time. This separation is powered by a quantum subroutine that appears throughout the literature on quantum speedups in learning, property testing, and computation: the multi-qubit SWAP operation [Harrow:2012gwf, Lloyd_2014, bădescu2017quantumstatecertification, Cincio_2018, montanaro2018surveyquantumpropertytesting, Huang_2022]. Here, we show that this super-exponential quantum speedup is completely degraded by only a single layer of noise.
@@
Our analysis proceeds by bounding the likelihood ratio between the leaf distributions generated by this learning tree when its measurement outcomes arise from the maximally mixed state versus from a fixed, Haar-random pure state. To do so, we focus on a fixed root-to-leaf path, utilizing the tree’s multilinear structure to reinterpret the learning problem. Under our reinterpretation, it suffices to bound the concentration of the output distributions of intermediate nodes in the tree, which we achieve by proving several technical lemmas regarding the noise channel and using the martingale formalism from [weiyuan_paulis]. Because each node concentrates by a factor exponentially small in the noise rate, every experiment is highly uninformative, and we obtain our lower bound.
@@
While we have shown that assessing the purity of a generic quantum state is fundamentally obstructed by noise, there are natural classes of “noise-robust” states for which purity testing remains feasible. A toy but instructive example arises in tensor-network models of the AdS/CFT correspondence [Maldacena_1999], such as the HaPPY code [happy_2015]. In essence, a HaPPY tensor network $T$ defines an isometric encoding $T:\mathcal{H}_{\text{bulk}}\to\mathcal{H}_{\text{bdy}}$ from a collection of “bulk” logical qudits on a finite hyperbolic lattice to “boundary” physical qudits arranged on a one-dimensional ring. In the AdS/CFT interpretation, $\mathcal{H}_{\text{bulk}}$ models a code subspace of low-energy quantum-gravitational degrees of freedom on a spatial slice of AdS, while $\mathcal{H}_{\text{bdy}}$ models the Hilbert space of the dual CFT on the boundary circle. The quantum error-correcting structure of $T$ discretely implements bulk-boundary reconstruction: local bulk operators can be reconstructed on many overlapping boundary regions, and the radial direction of the network plays the role of a renormalization group scale [Jahn_2021].
@@
This result should be viewed as a holographic counterpart of our noisy purity-testing lower bound. On the one hand, the HaPPY code converts local boundary erasures into highly suppressed logical errors on the bulk black hole degrees of freedom: as long as $R\gtrsim r+O(\log r)$ and the erasure rate is below threshold, a greedy decoder can approximately recover $\rho_{\mathrm{BH}}$ from the noisy boundary state with failure probability that is exponentially small in $4^{R-r}$. More broadly, it is believed (but not known) that HaPPY-type holographic codes may admit fully fault-tolerant realizations against local noise [numerics_happy_decoding_2020, Jahn_2021, Farrelly_2021]; in our setting we only appeal to their rigorously understood erasure-correction properties. Composing the decoding map with a two-copy SWAP test on the recovered bulk region therefore reproduces, up to small decoding errors, the ideal two-copy purity test and yields a constant-copy quantum protocol. On the other hand, any protocol restricted to single-copy measurements on the boundary reduces, via bulk-boundary isometry and data-processing, to single-copy purity testing on an $L_{r}$-qubit system, where $L_{r}=\Theta(4^{r-1})$ is the number of bulk legs in the excised region. Our general lower bound for noisy single-copy purity testing then implies a sample complexity of order $2^{\Theta(L_{r})}=2^{\exp(\Theta(r))}$, establishing the separation claimed in the theorem.
Our HaPPY code example illustrates the first of two strategies for useful noisy quantum learning, namely targeting physical systems with latent error-resilient properties.
2.3 Noise-dependent advantage in Pauli shadow tomography and metrology
@@
To further understand how noise reshapes quantum learning advantages, we now consider the well-studied problem of Pauli shadow tomography, namely estimating expectation values of (potentially mutually noncommuting) Pauli observables, which in ideal settings exhibits an exponential memory-sample tradeoff [chen2021exponentialseparationslearningquantum]. Any noiseless conventional protocol restricted to single-copy measurements, or to $k<n$ qubits of quantum memory, requires a number of samples exponential in nkn-k, whereas a quantum-enhanced learner with two-copy access (i.e. $n$ ancillary qubits of quantum memory) succeeds with only $O(n)$ samples using Bell-basis measurements; these two settings are depicted in Fig. 3(a) (memoryless or small-memory single-copy experiments) and Fig. 3(b) (architectures with a quantum memory register enabling multi-copy measurements). In Theorem 2.6, we refine this separation in the NBQP setting, quantifying the noise-dependence of lower bounds with and without ancillary quantum memory. Notably, even when $n$ qubits of memory are provided, enabling two-copy measurements, we show that order $(1-\lambda)^{-n}$ samples are necessary. We give a noisy 2-copy algorithm achieving this scaling up to a constant factor in the exponent, establishing a quantum advantage in noisy Pauli tomography that depends polynomially on the noise rate, and negating the exponential speedup of ideal Bell-measurement-based strategies.
@@
In the presence of constant depolarizing noise per qubit, any quantum algorithm without ancillary quantum memory which can identify a Pauli-structured state with high probability requires order $2^{n}f(\lambda)^{n}$ measurements, where $f(\lambda)\in[1,\infty)$ for $\lambda\in[0,1]$. Given an additional $k\leq n$ qubits of quantum memory, $\Omega(2^{n-k}(1-\lambda)^{-n})$ samples are still required. When $k=n$, there exists a quantum-enhanced learning algorithm with access to noisy two-copy measurements solving the Pauli-identification task using $\tilde{O}((1-\lambda)^{-4n})$ samples.
As an immediate corollary, we find a sample complexity separation between our two-copy algorithm and any single-copy strategy for the same Pauli identification task. This separation depends on the noise rate, and interpolates smoothly between the ideal exponential separation for Pauli tomography and a polynomial (in the noise rate) separation for $\Theta(1)$ local noise. A related separation is presented in Ref. [huang2022foundations]; that work considers diamond-norm gate noise and access to perfect state copies, a weaker noise model than NBQP, and gives a lower bound which considers only noiseless conventional experiments.
Corollary 2.7 (Noise-dependent quantum advantage in Pauli identification).
Let $N_{SC}$ be the optimal sample complexity for any single-copy algorithm for the $n$-qubit Pauli identification task from Theorem 2.6, and let $N_{TC}$ be the sample complexity of the two-copy algorithm given in Theorem 2.6. Then $N_{SC}=\Omega\big(N_{TC}^{a(\lambda)}\big)$, where $a(\lambda)$ is a function of noise rate such that when $\lambda=\Theta(1)$ and $n$ is large, $a(\lambda)=\Theta(\lambda^{-1})$, and as $\lambda\rightarrow 0$, $N_{SC}=\exp(\Omega(n))$ while $N_{TC}=\textnormal{poly}(\Theta(n))$.
@@
We next define learning trees for each experimental model. For models (i) and (ii), the output of each node of the tree is a classical bitstring. For both, we show that the distribution over bitstrings induced by the results of each experiment, given access to a $\mathds{1}+P$-type state, concentrate around the distribution resulting from experiments which instead query a maximally mixed state. Hence, every experiment is uninformative in distinguishing the two hypotheses, and exponentially many measurements are required to achieve an algorithm output distribution under one hypothesis that is far in total variation from the other (Theorem F.6 and Theorem F.8). The relationship between concentration of node-wise distributions and the final leaf output distribution is given by the martingale formalism developed in [weiyuan_paulis]. In model (ii), obtaining this bound requires quantifying the correlation between copies of the unknown state in terms of the size of the quantum memory and its lifetime, which we accomplish using a reformulation of the learning model in terms of Matrix Product States (Lemma F.11).
@@
For model (iii), nodes of the tree are joined by the state of the ancillary memory register after every measurement rather than a simple classical bitstring. Thus, we require a stronger approach: bounding the variation in entire root-to-leaf paths in the learning tree, following the approach from [chen2021exponentialseparationslearningquantum]. Using a probabilistic argument, we control the total variation of the leaf output distribution by bounding the number of paths which diverge substantially from the output distribution induced by the maximally mixed input state. To account for depolarizing noise in every copy of the state, we simplify the tensor-network analysis of [chen2021exponentialseparationslearningquantum] via a technical argument which provides a cleaner form of the action of our noise channel. With this approach, we find that the contribution of the noise to the sample complexity decouples from the number of memory qubits; hence, even given $n$ ancillary memory qubits, $\Omega((1-\lambda)^{-n/3})$ samples are still required in the noisy regime. We provide a two-copy algorithm which leverages noisy Bell sampling to match this asympotic lower bound up to constants in the exponent, and give a single-copy classical shadow algorithm for the broader Pauli shadow tomography which accounts for noise by using shallow-circuit unitary ensembles. Up to polynomial factors and constants in the exponents, our upper and lower bounds for the Pauli-identification task exhibit the same exponential dependence on $n$, $k$, and the noise rate $\lambda$, matching the known noiseless bounds from [chen2021exponentialseparationslearningquantum, weiyuan_paulis] as $\lambda\to 0$ and diverging to infinity as $\lambda\to 1$, where the task becomes information-theoretically impossible.
From the standpoint of near-term experiments, the most relevant regime is often not the asymptotic limit $n\to\infty$ at fixed noise. Instead, the instance size $n$ is effectively fixed by hardware, and the central question is how performance changes as other parameters, especially the noise rate $\lambda$, are improved. Our proof of Corollary 2.7 supports this viewpoint: for any fixed $n$, it gives an explicit $\lambda$-dependent separation between the optimal single-copy sample complexity $N_{SC}$ and the two-copy sample complexity $N_{TC}$, and thus provides a quantitative condition on $\lambda$ for obtaining a meaningful two-copy advantage. This perspective is consistent with the experimental results of Ref. [Huang_2022], which indicate that at modest instance sizes and sufficiently low noise, two-copy Pauli shadow protocols can substantially outperform single-copy baselines.
Our noise- and resource-aware analyses extend beyond state learning. Figure 4 contrasts ideal error-corrected metrology, which can sustain Heisenberg-limited scaling indefinitely, with the NBQP setting, where noisy probe-control couplings confine Heisenberg-limited performance to a finite sensing window before a crossover to the Standard Quantum Limit. In Theorem 2.8, we bound this precision-time tradeoff of Heisenberg-limited (HL) single-parameter metrology protocols under an effective noise model similar to NBQP.
Figure 4: (a) Ideal error-corrected metrology: interleaved quantum error correction (QEC) cycles protect an $N$-qubit probe during interrogation, sustaining Heisenberg-limited scaling in a noiseless sensing-control interface. (b) Natural error-corrected metrology: constant-rate faults at the sensing-control interface (red crosses, rate $R$) accumulate and restrict Heisenberg-limited performance to a finite sensing window (scaling to leading order as $\mathrm{poly}(N^{-1}R^{-1})$), beyond which the protocol crosses over to the Standard Quantum Limit.
Theorem 2.8 (A threshold for Heisenberg-limited metrology, informal).
Consider a single-parameter quantum sensing protocol in which a Hamiltonian $H(\omega)$ acts on an $N$-qubit probe, achieving Heisenberg-limited sensitivity when the sensing–control interface is noiseless. Under depolarizing noise of rate $R(\tau,\eta)$ (determined by the speed of control $\tau$ and native error strength $\eta$) at this interface, and without additional error mitigation, the protocol achieves the Heisenberg limit only up to a total evolution time inverse-polynomial in $NR$, beyond which it crosses over to the Standard Quantum Limit.
This result illustrates the value of noise-aware analysis beyond state preparation and measurement, namely that SPAM-robust metrology protocols can still fail under interstitial noise. It also highlights the connection between noisy quantum learning theory and practical metrology, demonstrating that even with access to quantum error correction, Heisenberg-limited metrological gains are delicate and tightly coupled to hardware constraints. The formal statement (Theorem G.2) expresses the threshold sensing time in terms of control speed, native noise strength, and quantum memory size. When all parameters except $N$ are $\Theta(1)$, HL sensitivity is essentially completely lost, as coherence-destroying errors accumulate on the same timescale as quantum control is executed in a single round. More generally, our bounds clarify the tradeoff between experimental resources and quantify how error rates must scale with $n$ to sustain an HL protocol over meaningful total time. For instance, an effective noise strength of $\eta\sim 1/N$ and $N\eta\tau\ll 1$ enables HL sensitivity for $O(1)$ time.
@@ 3 Outlook
We have studied quantum-enhanced experiments in the presence of noise through the lens of quantum learning theory. Our results show that many of the most striking idealized quantum advantages in learning, property testing, and metrology disappear, or become inapplicable, once interstitial noise and unprotected oracle systems are taken into account. At the same time, we identified settings where advantages survive, often in a weakened but still meaningful form, and clarified how these surviving advantages depend on noise strength, memory resources, and problem structure.
@@
Our work suggests several concrete directions for further development of noisy quantum learning theory. First, progress will require engaging more deeply with the physics of quantum many-body systems, rather than treating oracles as abstract. Many natural systems possess built-in robustness: renormalization-group structure and locality can protect long-range correlations [Kim_2017, Furuya_2022, Furuya_RGAQECC, goldman2024lindbladianexactrenormalizationdensity, Lake_2025], and thermalizing dynamics can generate noise-resilient macroscopic observables [Kastoryano_2025, Brandao_2019, Chesi_2010]. It will be important to identify and characterize classes of states, channels, and observables whose relevant features are intrinsically stable under realistic error models, and to turn such structure into quantitatively sharp, physically natural examples of noisy quantum advantage.
@@
Second, most known exponential quantum advantages in learning rest on highly entangled, global measurements, such as large-scale SWAP tests or Bell-basis measurements [Aharonov_Cotler_Qi_2022, huang2022foundations, chen2021exponentialseparationslearningquantum], that are simultaneously fragile to noise and misaligned with the local-control constraints of real devices which can couple to and manipulate experimental systems. This points toward a theory of noisy quantum learning under restricted resources, in which allowed operations may be shallow, geometrically local, few-qubit, or constrained to a small number of probe systems. Quantum probe tomography [chen2025quantumprobetomography] is an example of this philosophy, wherein a small number of probes interacting locally with a large system can still extract global information under noise. It will be important to understand systematically which restricted-access models admit robust (even if only polynomial) advantages over conventional experiments, and how those advantages trade off against architectural constraints.
@@
There is a close connection between noisy quantum learning and quantum sensing. Many sensing protocols can be viewed as learning problems about Hamiltonian parameters or state observables under stringent access and noise constraints [Zhou_2018, Huang_2023, Hu_2025]. Some ambitious proposals, such as quantum computation-enhanced sensing based on deep coherent control [Huang_2023, Hu_2025] or Grover-type oracles [allen2025quantumcomputingenhancedsensing], lose their asymptotic advantage in the NBQP noise model. This raises a quantitative question rather than a purely asymptotic one: how does the achievable quantum advantage degrade as a function of the noise rate, circuit depth, and available memory? Corollary 2.7, together with related work such as Ref. [huang2022foundations], suggest that in realistic noise regimes one should expect noise-dependent polynomial improvements rather than noise-independent exponential ones, but those polynomial advantages can still be practically meaningful.